OpenStudy (anonymous):

Allicin (C6H10OS2) contains 59.9 g of carbon, 21.7 g of hydrogen, 24.4 g of oxygen, and sulfur. What is the mass percent of sulfur in a 162 g sample of the compound

OpenStudy (sleepyjess):

@shrutipande9

OpenStudy (teddyiswatshecallsme):

C6H10OS2. mol mass = (C:12 x 6) + (H:1 x 10) + (O:16 x 1) + (S:32 x 2). = 72 + 10 + 16 + 64 = 162g/mol. Of this 'S' = (64 / 162) x 100% = 39.51 = mass percent of sample. (or 64g).